40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 642     Entry time: Mon Jul 7 16:30:08 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: PSL-PEM 16 days trend 
This morning the laser head temp was up to 20.3C because the laser chiller was overflowing.
I removed 700 cc water.

The PSL-FSS_RMTEMP became much more stable during the holidays as the psl enclosure was closed for 4 days

The high particle counts can be explaned by construction activity today.

The PMC & MZ PZT high voltages were out of range this morning.
Attachment 1: pempsl16d.jpg  125 kB  | Hide | Hide all
pempsl16d.jpg
ELOG V3.1.3-